On January 1, 2016, Drennen, Inc., issued $4.6 million face amount of 11-year, 10% stated rate bonds when market interest rates were 8%. The bonds pay semiannual interest each June 30 and December 31 and mature on December 31, 2026. Table 6-4, Table 6-5

a. Calculate the proceeds (issue price) of Drennen, Inc.'s, bonds on January 1, 2016, assuming that the bonds were sold to provide a market rate of return to the investor. (Round PV factor to 4 decimal places.)

  

b-1. Assume instead that the proceeds were $4,542,000. Use the horizontal model to record the payment of semiannual interest and the related discount amortization on June 30, 2016, assuming that the discount of $58,000 is amortized on a straight-line basis. (Use amounts with + for increases and amounts with – for decreases.)

b-2. Assume instead that the proceeds were $4,542,000. Record the journal entry to show the payment of semiannual interest and the related discount amortization on June 30, 2016, assuming that the discount of $58,000 is amortized on a straight-line basis. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

a. Bond issue price is the present value of future cash outflows discounted at market rate. = $5,264,751.31

Below are the workings for bond issue price calculation.

Refrence Drennen Inc. $46,00,000.00 5% Company Name Face Amount Bond Interest rate (Semin annually) Cash Interest Payment (Se

| | Calculation of Bond Issuance Price - WN 1 Interest Date Cash interest payment Discounting Factor market rate (4%) Present

b-1. Issue price is $4,542,000 and face value is $4,600,000 thus, bond is sold at discount, which means market rate should have been higher than issue price. Market rate is the rate at which interest expenses are discounted, Yield to Maturity will be the same as market rate.

Market Rate (Semi annual bond) = 2 x (1/2 x Coupon return + Pro-rated discount) / 1/2(Redemption price + purchase price)

A . 1/2 Coupon return = 1/2 (Coupon rate x FV) = $230,000

B. 1/2 Prorated discount = 1/2 (Redemption Price - Issue price ) / No. of periods = ($4,542,000 - $4,600,000)/ 22 periods = $2636.36

C. 1/2 Redemption price + purchase price) = 1/2 ($4,542,000 + $4,600,000) = $4,571,000

Market Rate = 2 x (A+B)/ C = 5.097% (Semi annually)

PS: Interest expense is always calculated by multiplying carrying value with market interest rate.

Carrying value (End of period) taal Bond Amortization table Cash Paid for Interest Interest Expense Amortization (Carrying vaTherefor Interest of $231,489 and discount amortization of $2636.36 is made on June 30, 2016

b-2. Bond discount is first debited to a payable account during issue of bond, later this amount is amortized in equal installments along the period of bond maturity.

I have included the journal entry for Jan 1, 2016 also, So that you get a perspective of which accounts are charged on June 30, 2016.

Debit Credit Jan 01, 2016 Account Cash A/C Bond discount amortizable To Bond payable (Being Bond issued) 45,42,000.00 58,000.